OpenStudy (anonymous):

Kevin and Randy Muise have a jar containing 53 coins, all of which are either quarters or nickels. The total value of the coins in the jar is $10.25. How many of each type of coin do they have? The jar contains ____ quarters. The jar contains ____ nickels.

OpenStudy (campbell_st):

you need to have a set of equations q + n = 53 ( the number of nickels and quarters) 0.25q + 0.1 n = 10.25 ( value of your coins) you need to solve them simultaneously

OpenStudy (anonymous):

how do you do that

OpenStudy (campbell_st):

ok... multiply the 2nd equation by 4 and subtract it from the 1st equation this will eliminate q and allow you to find the number of nickels q + n = 53 - q + 0.4 n = 41 ---------------- 0.6 n = 12 n = 12/0.6 so there are 20 nickels... how may quarters...?